# Energy-Based Model for Accurate Estimation of Shapley Values in Feature Attribution

### v3

@@ V-C Theoretical error bounds of EmSHAP and competitive methods
(EmSHAP estimation error upper bound) The expectation of the mean absolute deviation (MAD) between the estimation contribution function $\hat{v}(S)$ and the true contribution function $v(S)$ is upper bounded by $\frac{\sqrt{\pi}}{\sqrt{2K}}+\epsilon_{2}$, with $K$ being the number of sampling and $\epsilon_{2}$ a small positive number.
@@
$=\mathbb{E}\left[\left|\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right.\right.$
$\quad\left.\left.+\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]-\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})\right]\right|\right]$
@@
$\quad+\mathbb{E}\left[\left|\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]-\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})\right]\right|\right],$
@@
$\quad\mathbb{E}\left[\left|\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|\right]$ (35)
$=2\mathbb{E}\left[\left(\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right)_{+}\right]$
@@
$=\frac{\sqrt{\pi}}{\sqrt{2K}},$
[…]
$\quad\mathbb{E}\left[\left|\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]-\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})\right]\right|\right]$ (36)
$=\mathbb{E}\left[\left|\int f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d\mathbf{x}_{\bar{S}}\right.\right.$
$\quad\left.\left.-\int f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d\mathbf{x}_{\bar{S}}\right|\right]$
$=\mathbb{E}\left[\left|\int f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\frac{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d\mathbf{x}_{\bar{S}}\right.\right.$
$\left.\left.\quad-\int f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d\mathbf{x}_{\bar{S}}\right|\right]$
@@
By combining Eq.(35) and Eq.(36), the upper bound of $\mathbb{E}[|\hat{v}(S)-v(S)|]$ can be obtained, which is $\frac{\sqrt{\pi}}{\sqrt{2K}}+\epsilon_{2}$. ∎
@@
$\mathbb{E}[|\hat{v}(S)-v(S)|]$ $\leq 2(|D|+1)(|S|+1)\frac{\sqrt{\pi}}{\sqrt{2K}}||\bm{\Sigma}^{*^{-1}}||_{F}$ (39)
$\quad+2(|D|+1)^{\frac{5}{2}}(|S|+1)\frac{\sqrt{\pi}}{\sqrt{2K}}||\bm{\Sigma}^{*^{-1}}||^{2}_{F}+\epsilon_{3}^{2},$
where $\epsilon_{3}$ is a small positive number. $\bm{\Sigma}^{*}\in\mathbb{R}^{K\times K}$ is the expectation of Shapley kernel symmetric matrix, more detailed information about this matrix can be found in Ref. [12].
@@
The detailed proof can be found in Appendix. A. ∎
@@
(VAEAC estimation error upper bound) The expectation of the mean absolute deviation (MAD) between the contribution function $\hat{v}(S)$ estimated by VAEAC [17] and the true contribution function $v(S)$ is upper bounded by $\frac{\sqrt{\pi}}{\sqrt{2K}}+\epsilon_{2}+\delta$, where $K$ is the sampling number, $\epsilon_{2}$ is the probability density ratio between the estimation conditional probability $p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$ and the optimal conditional probability $p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$ estimated using VAE, $\delta$ is a positive number related to the amortization error of VAE.
$\small\left|\frac{p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}{p_{data}(\mathbf{x}_{\bar{S}}^{t}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}-1\right|\leq\delta.$ (40)
@@
The detailed proof can be found in Appendix. B. ∎
Theorems 2 and 3 analyzed the error bounds of two widely applied methods of KernelSHAP [12] and VAEAC [17]. Theorem 2 shows that as the sampling time $K$ increases, the first term of the right hand side of Eq.(39) vanishes, hence its error bound is governed by the second term. Since $||\bm{\Sigma}^{*^{-1}}||^{2}_{F}$ is a constant, it can be seen that the error bound of KernelSHAP is highly related to the dimension of input features $D$. That is to say, the estimation accuracy will decrease when the input feature dimension increases. On the other hand, comparing Theorem 1 and Theorem 3, it can be seen that the error bound of VAEAC is greater than the energy-based model as $\delta$ is always greater than zero. Hence it can be concluded that the proposed EmSHAP has a tighter error bound than KernelSHAP and VAEAC.
@@ V-D Validation of Error Bound through a toy example
Figure 3: Toy example for validating error bound of EmSHAP
@@ VI Case studies
In this section, feature attribution experiments are performed on four real-world tasks: image classification on MNIST [^3] MNIST dataset was acquired at http://yann.lecun.com/exdb/mnist/., question classification on TREC [^4] TREC dataset was acquired at https://cogcomp.seas.upenn.edu/Data/QA/QC/., tabular classification on adults’ income(ADTI) [^5] ADTI was acquired at https://archive.ics.uci.edu/dataset/2/adult., and time-series prediction of electric transformer’s oil temperature(ETT) [^6] ETT dataset was acquired at https://github.com/zhouhaoyi/ETDataset.. For the four tasks, EmSHAP is used to estimate Shapley values for feature attribution, which is compared with five competitive methods:
  • SamplingSHAP [7]: A Shapley value estimation method based on sampling from random permutation of features.
@@
  • KernelSHAP [12]: An effective method for calculating Shapley values which uses weighted least squares to calculate marginal Shapley values.
@@
Since the true Shapley value of the real-world datasets are difficult to calculate, a quantitative metric is used to evaluate the performance of feature attributions.
@@
Hyperparameter selection FastSHAP, VAEAC, and EmSHAP are based on neural networks, and their hyperparameter are summarized in Table II. All MLPs in the table include two residual connections. All models are trained using the Adam optimizer with a learning rate of 0.001. For the MNIST dataset, the batch size for all three models is set as 256. For the TREC and ADTI datasets, the batch size is set as 64, and the batch size of ETT is set to be 100. For EmSHAP, the minimum and maximum thresholds for dynamic masking are set as $\zeta_{\min}=0.2$ and $\zeta_{max}=0.8$, respectively, and the dimension of context vector $\gamma$ is set as 32. It is worth noting that increasing the latent dimensions may improve the accuracy of explanations, but it will also significantly increase computational cost. To achieve a trade-off between accuracy and computational efficiency, and to ensure a fair comparison, the hidden layer size is set as 32 for VAEAC and EmSHAP, and 128 for FastSHAP.
TABLE II: Network structure of feature attribution methods on real-world datasets
Method Network Prototype Latent
@@
VAEAC Full encoder MLP 32
@@
EmSHAP Proposal network GRU $2|D|+|\bm{\gamma}|$
@@
Table III summarizes the feature attribution performance and computation efficiency of different methods on four real-world datasets. All experiments are performed on a workstation equipped with an Intel Xeon Gold 6326 processor, 377G memory, and an RTX4080 graphics processing unit. The calculations are performed using Python 3.9 and PyTorch 2.4.1. The calculation time and memory usage for all methods are also compared and listed in Table III. From Table III it can be seen that EmSHAP yields the highest SIC AUC-ADD and the lowest SIC AUC-DEL in all cases, indicating that EmSHAP has the best feature attribution performance. For computation efficiency, Table III shows that methods that do not involve training stage are more efficient. However, for methods that involving both training and inference stages, it can be seen that EmSHAP has the highest efficiency as it requires the least amount of time.
TABLE III: Performance comparison of feature attribution for competitive methods on real-world datasets
Dataset Explainer SIC AUC Run time Memory usage
@@
MNIST SamplingSHAP 0.893 0.269 /// 2.55s /// $\mathbf{1464.81}$MB
SVARM 0.257 0.125 /// 2.03s /// 1647.98MB
KernelSHAP 0.861 0.134 /// 1.84s /// 1714.42MB
FastSHAP 0.861 0.205 3302.66s $\mathbf{0.19}$s 2900.80MB 1958.56MB
@@
EmSHAP $\mathbf{0.965}$ $\mathbf{0.037}$ $\mathbf{2977.10}$s 1.63s $\mathbf{1823.91}$MB 1879.68MB
TREC SamplingSHAP 0.888 0.155 /// 1.62s /// 978.40MB
SVARM 0.846 0.187 /// 1.23s /// 980.22MB
KernelSHAP 0.875 0.135 /// 0.34s /// 993.72MB
FastSHAP 0.815 0.194 204.75s $\mathbf{0.14}$s 1168.51MB $\mathbf{890.36}$MB
VAEAC 0.725 0.316 646.19s 3.10s 1091.89MB 1071.34MB
EmSHAP $\mathbf{0.948}$ $\mathbf{0.103}$ $\mathbf{162.75}$s 0.19s $\mathbf{1085.62}$MB 994.82MB
ADTI SamplingSHAP 0.925 0.494 /// 0.20s /// 718.47MB
SVARM 0.927 0.518 /// $\mathbf{0.18s}$ /// 696.97MB
KernelSHAP 0.913 0.465 /// 1.43s /// 703.79MB
FastSHAP 0.914 0.535 295.87s 0.26s $\mathbf{717.17}$MB $\mathbf{696.64}$MB
@@
EmSHAP $\mathbf{0.933}$ $\mathbf{0.411}$ $\mathbf{270.90}$s 0.87s 966.58MB 946.61MB
ETT SamplingSHAP 0.893 0.526 /// 0.08s /// 761.11MB
SVARM 0.875 0.342 /// 0.05s /// 758.97MB
KernelSHAP 0.907 0.289 /// 7E-3s /// $\mathbf{534.54}$MB
FastSHAP 0.863 0.402 19.39s $\mathbf{1E-3}$s 707.49MB 586.20MB
VAEAC 0.902 0.359 46.91s 0.33s $\mathbf{530.85}$MB 568.14MB
EmSHAP $\mathbf{0.915}$ $\mathbf{0.243}$ $\mathbf{16.55}$s 4E-3s 554.11MB 540.46MB

### v4

@@ V-C Theoretical error bounds of EmSHAP and competitive methods
(EmSHAP estimation error upper bound) The expectation of the mean absolute deviation (MAD) between the estimation contribution function $\hat{v}(S)$ and the true contribution function $v(S)$ is upper bounded by
$\small\mathbb{E}[|\hat{v}(S)-v(S)|]\leq\frac{\sqrt{\pi}}{\sqrt{2K}}+\epsilon_{2},$ (31)
where $K$ is the number of sampling and $\epsilon_{2}$ is a small positive number.
@@
[…]
$\qquad\left.\left.+\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]-\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})\right]\right|\right]$
@@
$\qquad+\mathbb{E}\left[\left|\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]-\mathbb{E}_{p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})\right]\right|\right],$
@@
$\qquad\mathbb{E}\left[\left|\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right|\right]$
$=2\mathbb{E}\left[\left(\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\right)_{+}\right]$ (33)
$=2\int_{0}^{+\infty}\mathbb{P}\left(\frac{1}{K}\sum_{k=1}^{K}f(\mathbf{x}_{\bar{S}}^{(k)},\mathbf{x}_{S}^{t})-\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]\geq\epsilon_{3}\right)d\epsilon_{3}$
$\leq 2\int_{0}^{+\infty}\exp(-2K\epsilon_{3}^{2})d\epsilon_{3}=\frac{\sqrt{\pi}}{\sqrt{2K}},$
where $(\cdot)_{+}$ is $max(\cdot,0)$, the integral in Eq.(V-C) uses the equation $\mathbb{E}[u]=\int_{0}^{+\infty}\mathbb{P}(u>\epsilon)d\epsilon$ for a non-negative random variable $u$, the inequality in Eq.(V-C) holds according to Corollary 2. Hence, the statistical error is bounded by $\frac{\sqrt{\pi}}{\sqrt{2K}}$. The derivation of the approximation error is given as follows.
$\qquad\mathbb{E}\left[\left|\mathbb{E}_{p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})\right]-\mathbb{E}_{p_{data}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})}\left[f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})\right]\right|\right]$ (34)
@@
$\qquad\left.\left.-\int f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d\mathbf{x}_{\bar{S}}\right|\right]$
@@
$\left.\left.\qquad-\int f(\mathbf{x}_{\bar{S}}^{t},\mathbf{x}_{S}^{t})p_{data}(\mathbf{x}^{t}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d\mathbf{x}_{\bar{S}}\right|\right]$
@@
By combining Eq.(V-C) and Eq.(34), the upper bound of $\mathbb{E}[|\hat{v}(S)-v(S)|]$ can be obtained, which is $\frac{\sqrt{\pi}}{\sqrt{2K}}+\epsilon_{2}$. ∎
@@
[…]
where $\epsilon_{3}$ is a small positive number related to $\epsilon_{1}$ and $\epsilon_{1}^{\prime}$. $\bm{\Sigma}^{*}\in\mathbb{R}^{K\times K}$ is the expectation of Shapley kernel symmetric matrix, more detailed information about this matrix can be found in Ref. [14].
@@
The detailed proof can be found in Appendix A. ∎
@@
(VAEAC estimation error upper bound) The expectation of the mean absolute deviation (MAD) between the contribution function $\hat{v}(S)$ estimated by VAEAC [18] and the true contribution function $v(S)$ is upper bounded by,
$\small\mathbb{E}[|\hat{v}(S)-v(S)|]\leq\frac{\sqrt{\pi}}{\sqrt{2K}}+\epsilon_{2}+\delta,$ (38)
where $K$ is the sampling number, $\epsilon_{2}$ is the probability density ratio between the estimation conditional probability $p_{\theta}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$ and the optimal conditional probability $p_{\theta^{*}}(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$ estimated using VAE, $\delta$ is a positive number related to the amortization error of VAE.
@@
The detailed proof can be found in Appendix B. ∎
Theorems 2 and 3 analyzed the error bounds of two widely applied methods of KernelSHAP [14] and VAEAC [18]. Theorem 2 shows that as the sampling time $K$ increases, the first term of the right hand side of Eq.(37) vanishes, hence its error bound is governed by the second term. Since $\bm{\Sigma}^{*^{-1}}||^{2}_{F}$ is a constant, it can be seen that the error bound of KernelSHAP is highly related to the dimension of input features $D$. That is to say, the estimation accuracy will decrease when the input feature dimension increases. On the other hand, comparing Theorem 1 and Theorem 3, it can be seen that the error bound of VAEAC is greater than the energy-based model as $\delta$ is always greater than zero. Hence it can be concluded that the proposed EmSHAP has a tighter error bound than KernelSHAP and VAEAC.
Remark 2.
The above analysis differs from existing work [56, 14, 57, 30]. Previous methods primarily address marginal Shapley values, where errors stem from sampling or regression approximations. Moreover, they are often tailored to specific scenarios like correlated features [58], distribution shifts [59], or non-parametric settings [60]. While these are valuable, they remain specialized. In contrast, Theorem 1 establishes an error bound for conditional Shapley values within an explicit generative modelling framework. This approach is more general in two key aspects: i) it applies to any generative model-based conditional estimator (e.g., VAEAC), and ii) when the conditional distribution is accurately modelled, the bound resembles the concentration inequalities-based bounds of marginal methods (e.g., KernelSHAP). Thus, our analysis provides a unifying framework where sampling- and regression-based approaches emerge as special cases under simplified assumptions.
@@ V-D Validation of Error Bound through a toy example
Figure 4: Toy example for validating error bound of EmSHAP.
@@ VI Case studies
In this section, feature attribution experiments are performed on four real-world tasks: image classification on MNIST [^3] MNIST dataset was acquired at http://yann.lecun.com/exdb/mnist/., tabular classification on adults’ income(ADTI) [^4] ADTI was acquired at https://archive.ics.uci.edu/dataset/2/adult., time-series prediction of electric transformer’s oil temperature(ETT) [^5] ETT dataset was acquired at https://github.com/zhouhaoyi/ETDataset., and automated medical coding on MIMIC-III [^6] MIMIC-III dataset was acquired at https://mimic.physionet.org/.. For the four tasks, EmSHAP is used to estimate Shapley values for feature attribution and is compared with seven competitive methods:
  • SamplingSHAP [7]: A Shapley value estimation method based on sampling from a random permutation of features.
@@
  • SHAP-IQ [38]: A sampling-based approximation method for calculating Shapley interaction quantification.
  • KernelSHAP [14]: An effective method for calculating Shapley values, which uses weighted least squares to calculate marginal Shapley values.
  • ST-SHAP[30]: A regression-based method utilizing neighbor selection to achieve a more stable version of KernelSHAP.
@@
Since the true Shapley value of the real-world datasets is difficult to calculate, a quantitative metric is used to evaluate the performance of feature attributions.
@@
Hyperparameter selection Before model training, hyperparameters should be determined. For SamplingSHAP, SVARM, and SHAP-IQ, the maximum sampling number (also called evaluation budget) is set as 1000, while for KernelSHAP and ST-SHAP, the number of sampling is fixed at 1000. FastSHAP, VAEAC, and EmSHAP are neural network-based methods, and their hyperparameters are summarized in Table I. In Table I, all MLPs include two residual connections, and the models are trained with the Adam optimizer under a learning rate of 0.001 for 100 epochs, with batch sizes of 256 for MNIST, 64 for ADTI, 100 for ETT, and 16 for MIMIC-III. For VAEAC and EmSHAP, the number of sampling from the conditional distribution is set as 20, and in EmSHAP, the number of importance sampling from the proposal distribution is also set as 20 following Ref. [64]. The thresholds for dynamic masking vary from $\zeta_{\min}=0.2$ to $\zeta_{\max}=0.8$, and the context vector dimension $\gamma$ is fixed at 32. To balance between accuracy and computational cost, the hidden layer size is set to be 32 for VAEAC and EmSHAP, and 128 for FastSHAP. The architectures of the predictive models to be explained in different case studies are provided in Appendix C.
TABLE I: Network structure of feature attribution methods on real-world datasets
Method
Network
Prototype
Latent
@@
VAEAC
Full encoder MLP 32
@@
EmSHAP
Proposal network GRU $2|D|+|\bm{\gamma}|$
@@
Table II summarizes the feature attribution performance and computation efficiency of different methods on four real-world datasets. All experiments are performed on a workstation equipped with an Intel Xeon Gold 6326 processor, 377 GB of memory, and an RTX4080 graphics processing unit. The calculations are performed using Python 3.9 and PyTorch 2.4.1. The calculation time and memory usage for all methods are also compared and listed in Table II. Table II shows that EmSHAP consistently achieves the highest SIC AUC-ADD in all datasets, and the lowest SIC AUC-DEL across most datasets, indicating superior feature attribution performance. Regarding computational efficiency, methods without a training stage generally exhibit higher efficiency [^7] Our version can be accessed on https://github.com/icanreachyou/EmSHAP..
TABLE II: Performance comparison of feature attribution for competitive methods on real-world datasets
Dataset
Explainer
SIC AUC Run time Memory usage
@@
MNIST
[…]
FastSHAP 0.861 0.205 3302.66s 0.19s 2900.80MB 1958.56MB
@@
EmSHAP 0.965 0.037 2977.10s 1.63s 1823.91MB 1879.68MB
ADTI
[…]
FastSHAP 0.914 0.535 295.87s 0.26s 717.17MB 696.64MB
@@
EmSHAP 0.933 0.411 270.90s 0.87s 966.58MB 946.61MB
ETT
[…]
FastSHAP 0.863 0.402 19.39s 1E-3s 707.49MB 586.20MB
VAEAC 0.902 0.359 46.91s 0.33s 530.85MB 568.14MB
EmSHAP 0.915 0.243 16.55s 4E-3s 554.11MB 540.46MB
MIMIC
[…]
FastSHAP 0.617 0.424 780.72s 1.17s 1254.49MB 987.54MB
VAEAC 0.572 0.466 13262.06s 128.65s 2762.11MB 2769.82MB
EmSHAP 0.699 0.384 1168.72s 59.93s 1165.51MB 1168.10MB